# Spin noise spectroscopy of an alignment-based atomic magnetometer

### v1

@@ III.2.1 Varying the noise spectral density
Figures 4(a) and 4(b) show the variation of the amplitude and line-width, respectively, for each of the Lorentzian peaks fitted around $f=0$, $f_{L}$ and $2f_{L}$, as a function of the white-noise strength $V_{\mathrm{noise}}$ in $\mathrm{mV^{2}_{rms}}$ units (same colouring as in the main plot of Fig. 3). The data was collected with a fixed, $\theta=25^{\circ}$, polarisation angle of the input beam. As described above, for low white-noise strengths, $V_{\mathrm{noise}}<1~{}\mathrm{V_{rms}}$, the fitting procedure was done while treating each of the peaks separately, whereas for higher values of noise a single fitting-function was used.
@@
Crucially, the experimental results are consistent with the theory, in particular, the amplitudes of the three (absorptive) peaks follow the functional dependences predicted by Eq. (III.1), with e.g. quadratic and linear dependences: $\widetilde{p}_{0}^{a}~{}\propto~{}f_{\mathrm{noise}}^{2}$, $\widetilde{p}_{1}^{a}\propto f_{\mathrm{noise}}$, and $\widetilde{p}_{2}^{a}\propto f_{\mathrm{noise}}^{2}$ with $f_{\mathrm{noise}}\propto V^{2}_{\mathrm{noise}}$ (now in Hz); easily verifiable from Fig. 4(a) in the low white-noise regime. On the other hand, the corresponding line-widths depicted in Fig. 4(b) follow quadratic dependences in $V_{\mathrm{noise}}$ predicted by Eq. (37), which allows us to determine the effective dissipation rates (23) that take values $\widetilde{\Gamma}_{0}=27(3)$ Hz, $\widetilde{\Gamma}_{1}=33(3)$ Hz and $\widetilde{\Gamma}_{2}=30(3)$ Hz for $f=0$, $f_{L}$ and $2f_{L}$ peaks, respectively. Moreover, in both cases we obtain the proportionality constant in $f_{\mathrm{noise}}=cV^{2}_{\mathrm{noise}}$ to be best approximated by $c=1.33(2)\times 10^{-5}~{}\mathrm{Hz/mV^{2}_{rms}}$, which is in agreement with the value obtained independently via the calibration described in App. D. Although the theory does not allow us to predict the proportionality constant $C$ shared by all the peak amplitudes in Eq. (III.1), we determine its value separately for each $\widetilde{p}^{a}_{0}$, $\widetilde{p}^{a}_{1}$ and $\widetilde{p}^{a}_{2}$ in Fig. 4(a). We obtain $C_{0}=2.97(7)\times 10^{-6}~{}\mathrm{V}^{2}$, $C_{1}=1.88(4)\times 10^{-6}~{}\mathrm{V}^{2}$ and $C_{3}=1.18(3)\times 10^{-6}~{}\mathrm{V}^{2}$, respectively, which consistently are of the same order.
Last but not least, let us comment that our theoretical model is incapable of predicting shifts in the central frequencies of the peaks, as it would force $f_{j}=j\,f_{L}$ in Eq. (42) with the corrections to $\omega_{j}=2\pi f_{j}$ predicted by Eq. (36) not only being negative, but also appearing at a negligible $O(\omega_{\mathrm{noise}}^{2}/\Omega_{L}=f_{\mathrm{noise}}^{2}/f_{L})$ order—ranging between $10^{-5}$ Hz and $1$ Hz for parameters of our experiment. Nonetheless, we investigate explicitly the shifts of central frequencies $f_{1}\approx f_{L}$ and $f_{2}\approx 2f_{L}$ with the white-noise strengths in Fig. 5. Surprisingly, these increase linearly with $f_{\mathrm{noise}}$ (quadratically with $V_{\mathrm{noise}}$) reaching up to $\Delta f_{j}\coloneqq f_{j}-jf_{L}\approx 1$ kHz for the strongest white-noise applied. Although the stochastic dynamics (24) does not anticipate such a phenomenon, it can be incorporated phenomenologically by letting the “effective” Larmor frequencies depend on the white-noise strength, so that assuming the shifts to depend linearly on the noise density we obtain $\Delta f_{j}=\chi_{j}f_{\mathrm{noise}}$ with $\chi_{1}=4.0$ and $\chi_{2}=6.2$ for the data of Fig. 5.
Figure 5: Central frequency of the PSD peaks as a function of the noise spectral density ($f_{\mathrm{noise}}$), which is varied from 7 $\mathrm{mV_{rms}}$ to 3500 $\mathrm{mV_{rms}}$ with the light polarisation angle being fixed to $\theta=25^{\circ}$. Although the shifts observed are not predicted by our theoretical model, we fit the linear (quadratic) dependence of $\Delta f_{j}\coloneqq f_{j}-jf_{l}$ on $f_{\mathrm{noise}}$ ($V_{\mathrm{noise}}$) to obtain phenomenological shift dependencies $\Delta f_{j}=\chi_{j}f_{\mathrm{noise}}$ with $\chi_{1}=4.0$ and $\chi_{2}=6.2$ for the peaks centred around the frequencies $f_{L}$ ($j=1$, green solid) and $2f_{L}$ ($j=2$, blue dashed).
@@ III.2.2 Varying the light polarisation angle
Finally, as the angular dependence of the peak amplitudes (III.1) separates from all other parameters in the form of functions $h[\theta]$ and $g[\theta]$ stated in Eqs. (40) and (41), we verify this explicitly in the low white-noise regime with $V_{\mathrm{noise}}=$140 $\mathrm{mV_{rms}}$ (each peak fitted separately according to Eq. (42)), by varying the polarisation angle of the incoming light from $\theta=-20^{\circ}$ to $\theta=120^{\circ}$. The results are presented in Fig. 6 with measurements reproducing almost exactly the predicted angular behaviours. As the line-widths of the peaks do not vary significantly when varying $\theta$ (data not shown), we may use the values $\widetilde{\gamma}_{j}$ determined for this low value of noise in Fig. 4(a) and compute separately the exact proportionality constants for each $\widetilde{p}^{a}_{j}$ such that the functions $h[\theta]$ and $g[\theta]$ are most accurately reproduced. In this way, while accounting also for a common angular offset $\theta\to\theta+\delta\theta$ with $\delta\theta=0.77(7)^{\circ}$ in our setup, we obtain $C_{0}=3.1(5)\times 10^{-6}$ V, $C_{1}=2.0(1)\times 10^{-6}$ V and $C_{2}=1.3(4)\times 10^{-6}$ V, which are consistently of similar magnitude and almost the same as the ones determined above when varying the white-noise strength.
Figure 6: Amplitudes of the PSD peaks as a function of the light polarisation angle, which is varied from $\theta=-20^{\circ}$ to $\theta=120^{\circ}$ with the white-noise strength set to $V_{\mathrm{noise}}=140~{}\mathrm{mV_{rms}}$. The fits predicted theoretically by Eq. (III.1), in particular, the angular dependences $h[\theta]$ and $g[\theta]$ stated in Eqs. (40) and (41), are shown as solid curves for the peaks centred around frequencies 00 (red), $f_{L}$ (green) and $2f_{L}$ (blue).
@@ IV Conclusions
On one hand, the added noise amplifies the Larmor-induced peaks to be clearly visible within the power spectral density above the level of detection noise, so that thanks to our model the device can be used as a scalar magnetometer to track fast changes (beyond the magnetometer bandwidth [46]) of the strong magnetic field in real time [59, 60]. On the other, the induced noise can be used to naturally perform sensing tasks in the so-called covert manner [61], i.e. so that an adversary having access to the output of the magnetometer would not be able to recover the signal without possessing the pre-calibrated dynamical model that we propose. In this sense, we expect our model to be useful also for tracking time-varying signals encoded in oscillating RF-fields (amplitude/phase) directed perpendicularly to both the scalar field and the added noise.
@@ ACKNOWLEDGEMENTS
This work was supported by the QuantERA grant C’MON-QSENS! funded by the Engineering and Physical Sciences Research Council (EPSRC) (Grant No. EP/T027126/1). Project C’MON-QSENS! is also supported by the National Science Centre (2019/32/Z/ST2/00026), Poland under QuantERA, which has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement no. 731473. The work was also supported by the Novo Nordisk Foundation (Grant No. NNF20OC0064182) and the UK Quantum Technology Hub in Sensing and Timing, funded EPSRC (Grant No. EP/T001046/1).
@@ Appendix A Evolution in the basis of spherical tensor operators
The evolution of the atomic ensemble without the effect of pumping, described by Eq. (19), is written in terms of the density matrix (1). In this appendix we show how to write it in terms of the vector of all spherical tensor components $\textbf{m}_{t}^{\mathrm{full}}$. In particular, since the right-hand side of the equation is linear in terms of $\rho$, we would like to find specific linear operators that act on $\textbf{m}_{t}^{\mathrm{full}}$ that correspond to specific operations performed on $\rho$, and enable us to write the dynamical equation in the form of (21). We also show that, since the resulting operators do not couple spherical tensor components of different rank, we only need rank-2 component for a full description of the problem under consideration.
@@ A.1 Rotations induced by magnetic fields
Spherical-tensor operators, ${{\mathcal{T}}}^{(\kappa)}_{q}$, form a convenient basis for density matrices of a system with fixed angular momentum $F$:
@@
Using Eq. (51) we could analogously find, that commuting the density matrix with $\hat{F}_{x}$ and $\hat{F}_{y}$ is equivalent to acting the operators $\textbf{J}^{(\kappa)}_{x}$ and $\textbf{J}^{(\kappa)}_{y}$, respectively, on $\textbf{m}_{t}^{(\kappa)}$. We use this to also obtain:
$\textbf{A}_{\mathrm{noise}}=-\mathrm{i}\sqrt{\omega_{\mathrm{noise}}}\textbf{J}^{(2)}_{x}$ (53)
@@ B.1 Effective Lorentzian form
where $\tilde{k}_{\alpha}$, $\tilde{k}_{\beta}$ are the components of the k vector and $\tilde{Q}_{\alpha\beta}$ are the components of the matrix $\textbf{G}\,\bm{\sigma}\,\textbf{G}^{\mathrm{T}}$, in both cases written in the eigenbasis of F. As one can see, the resulting spectrum is a sum of symmetric (absorptive) and antisymmetric (dispersive) Lorentzian peaks, whose widths are the opposites of real parts ($\gamma_{\alpha}$), and central frequencies are the imaginary parts of the eigenvalues and their opposites (±$\pm\omega_{\alpha}$):
@@ Appendix C Analytic prediction of peak amplitudes
Figure 7: Amplitude calibration for the noise spectral density ($\omega_{\mathrm{noise}}$) (a) Fourier Transform of 100 averages of 0.01 seconds of data. The data is sampled at a rate of 40MHz and we have a 1MHz low pass filter. (b) Frequency averages of the data for varying white noise amplitude fitted to $f_{\textrm{noise}}=cV_{\mathrm{noise}}^{2}=\frac{\omega_{\mathrm{noise}}}{2\pi}$.
@@
because we need the second order of expansion in $\frac{1}{\Omega_{L}}$. Using this approximation we obtain Eqs. (III.1) from the main text, and:
@@ C.1 Irrelevance of the dispersive contributions
We note that we have not verified the predicted relations for the dispersive contributions of the line shape, see Eq. (91) and Eq. (92) in App. C. This is due to their contribution being below the noise floor of our system. We can verify that this is as expected for a Larmor frequency of approximately 10 kHz by calculating the ratios of the peak height equations and using the linewidths found in the experiment, $\Gamma_{1}=33$ Hz and $\Gamma_{2}=30$ Hz. We find $p^{d}_{1}/p^{a}_{1}$ from Eq. (91) divided by Eq. (38b) and we find $p^{d}_{2}/p^{a}_{2}$ from Eq. (92) divided by Eq. (38c). The calculations were done for the lowest white noise amplitude of $7~{}\mathrm{mV_{rms}}$ and the highest white noise amplitude of $3.5~{}\mathrm{V_{rms}}$. For $\mathrm{V_{noise}}=7~{}\mathrm{mV_{rms}}$ we find $p^{d}_{1}/p^{a}_{1}=1.0\times 10^{-4}$ and $p^{d}_{2}/p^{a}_{2}=3.0\times 10^{-4}$. Furthermore, when $\mathrm{V_{noise}}=3.5~{}\mathrm{V_{rms}}$ we find $p^{d}_{1}/p^{a}_{1}=2.9\times 10^{-5}$ and $p^{d}_{2}/p^{a}_{2}=1.6\times 10^{-4}$. Hence, the dispersive contribution is at least 4 orders of magnitude smaller than the absorptive contribution. Hence, this is significantly below our noise floor for both peaks, as seen experimentally.
Appendix D Amplitude calibration for the noise spectral density $\omega_{\mathrm{noise}}$
In order to determine an experimental value for the constant $\omega_{\mathrm{noise}}=2\pi f_{\mathrm{noise}}=2\pi cV_{\mathrm{noise}}^{2}$ we recorded 100 0.01 second time traces of the white noise generator at varying noise amplitudes. A 1 MHz low pass filter was used to prevent the effect of aliasing. The data was then Fourier transformed to find the power spectral density. Using the power spectral density together with the coil calibration factor (10.1 $\mathrm{nT/mV_{rms}}$) and the gyromagnetic ratio of Caesium (3.5 Hz/nT) we can determine the PSD in the units of $\mathrm{Hz}^{2}/\mathrm{Hz}$. The results are shown in Fig. 7.

### v2

@@ III.2.1 Varying the noise spectral density
Figures 4(a) and 4(b) show the variation of the amplitude and linewidth, respectively, for each of the Lorentzian peaks fitted around $f=0$, $f_{L}$ and $2f_{L}$, as a function of the noise spectral density $f_{\mathrm{noise}}$ (in Hz). The data was collected with a fixed, $\theta=25^{\circ}$, polarisation angle of the input beam. As described above, the fitting procedure was done using a single fitting-function. The applied voltage to generate the noisy magnetic field was varied from 7 $\mathrm{mV_{rms}}$ to 3500 $\mathrm{mV_{rms}}$. To convert this to spectral density (in Hz) we determine the proportionality constant in $f_{\mathrm{noise}}=c\mathrm{V^{2}_{noise}}$ as $c=1.33(3)\times 10^{-5}~{}\mathrm{Hz/mV^{2}_{rms}}$, which is possible via the calibration procedure described in App. D.
@@
Crucially, the experimental results are consistent with the theory, in particular, the amplitudes of the three (absorptive) peaks follow the functional dependences predicted by Eq. (39), with e.g. quadratic and linear dependences: $\widetilde{p}_{0}^{a}~{}\propto~{}f_{\mathrm{noise}}^{2}$, $\widetilde{p}_{1}^{a}\propto f_{\mathrm{noise}}$, and $\widetilde{p}_{2}^{a}\propto f_{\mathrm{noise}}^{2}$; easily verifiable from Fig. 4(a) in the low white-noise regime. Performing the full fitting procedure, we allow the proportionality constants in Eq. (39) to differ ($C_{j}$ for each $p^{a}_{j}$), while determining common dissipation rates that are most consistent with the data ($\Gamma_{0}$, $\Gamma_{1}$ and $\Gamma_{2}$). We obtain $C_{0}=2.1(2)\times 10^{-6}~{}\mathrm{V}^{2}$, $C_{1}=3.4(6)\times 10^{-6}~{}\mathrm{V}^{2}$ and $C_{2}=1.0(1)\times 10^{-6}~{}\mathrm{V}^{2}$ for each peak, respectively, which consistently are of the same order, whereas the best-fitted dissipation rates read $\widetilde{\Gamma}_{0}=29(1)$ Hz, $\widetilde{\Gamma}_{1}=43(4)$ Hz and $\widetilde{\Gamma}_{2}=29(2)$ Hz. On the other hand, the corresponding linewidths depicted in Fig. 4(b) follow linear dependences in $f_{\mathrm{noise}}$ with the slope almost identical to the proportionality constants predicted by Eq. (38). Moreover, their offsets at $f_{\mathrm{noise}}$ allows us to independently determine the dissipation rates (23) as $\widetilde{\Gamma}_{0}=31(3)$ Hz, $\widetilde{\Gamma}_{1}=34(1)$ Hz and $\widetilde{\Gamma}_{2}=34(3)$ Hz. These are in a good agreement (within $\approx\!10$Hz) with the ones found when fitting the three peak-amplitudes in Fig. 4(a).
Last but not least, let us comment on the ability of our theoretical models to predict changes in the central frequencies of the peaks. Fig. 5 shows how the central frequency, for the peak around $f_{L}$ (green dots) and $2f_{L}$ (blue squares), vary as a function of the noise spectral density in the experiment. The theoretically predicted relation, Eq. (37), is then fit to the experimental data (dashed lines) with $f_{L}=\Omega_{L}/(2\pi)=9435(1)$ Hz, being the only free parameter. It can be seen that the theory agrees reasonably well with the experimental data with the central frequency slightly decreasing in the high noise regime for the peak at $f_{L}$ and remaining almost the same for the second peak. Note that the error on the values for the central frequencies are within ±$\pm 5$ Hz. These errors come from experimental imperfections, however, such as the central frequency drifting in the time it takes for all of the data to be taken, and errors in fitting the data. It is noted here that the errors in fitting the central frequency at $2f_{L}$ will be larger than at $f_{L}$ due to the peaks being at least an order of magnitude smaller and hence closer to the noise floor of the experiment. Bearing in mind all the aforementioned imperfections, one may conclude that theory predicts well the dependence of the central frequencies on the noise density for both peaks.
Figure 5: Central frequency of the PSD peaks as a function of the noise spectral density ($f_{\mathrm{noise}}$), which is varied from $6.53\times 10^{-4}$ Hz to 163 Hz with the light polarisation angle being fixed to $\theta=25^{\circ}$. The theoretical predictions of the change in central frequency (Eq. (37)) are shown over the full range of the noise spectral densities applied (dashed lines) with $f_{L}=9435(1)$ Hz fitted to the experimental data for $f_{1}\approx f_{L}$ (green dots) and $f_{2}\approx 2f_{L}$ (blue squares). The overall error in the experimentally determined central frequencies is appr. ±$\pm 5$ Hz, arising both from experimental imperfections (e.g. drifts, background-noise subtraction) and the fitting procedure.
@@ III.2.2 Varying the light polarisation angle
Finally, as the angular dependence of the peak amplitudes (39) separates from all other parameters in the form of functions $h[\theta]$ and $g[\theta]$ stated in Eqs. (41) and (42), we verify this explicitly in the low white-noise regime with $V_{\mathrm{noise}}=$140 $\mathrm{mV_{rms}}$, corresponding to $f_{\mathrm{noise}}=0.26~{}\mathrm{Hz}$, by varying the polarisation angle of the incoming light from $\theta=-20^{\circ}$ to $\theta=120^{\circ}$. It is noted that the Larmor frequency used for varying the polarisation angle was $\sim 9.6$ kHz. The results are presented in Fig. 6 with measurements reproducing almost exactly the predicted angular behaviours. As the linewidths of the peaks do not vary significantly when varying $\theta$ (data not shown), we may use the values $\widetilde{\gamma}_{j}$ determined for this low value of noise in Fig. 4(a) and compute separately the exact proportionality constants for each $\widetilde{p}^{a}_{j}$ such that the functions $h[\theta]$ and $g[\theta]$ are most accurately reproduced. In this way, while accounting also for a common angular offset $\theta\to\theta+\delta\theta$ with $\delta\theta=0.77(7)^{\circ}$ in our setup, we obtain $C_{0}=3.9(6)\times 10^{-6}$ V, $C_{1}=2.1(1)\times 10^{-6}$ V and $C_{2}=1.5(4)\times 10^{-6}$ V, which are consistently of similar magnitude and almost the same as the ones determined above when varying the white-noise strength.
Figure 6: Amplitudes of the PSD peaks as a function of the light polarisation angle, which is varied from $\theta=-20^{\circ}$ to $\theta=120^{\circ}$ with the white-noise strength set to $V_{\mathrm{noise}}=140~{}\mathrm{mV_{rms}}$, corresponding to $f_{\mathrm{noise}}=0.26~{}\mathrm{Hz}$. The fits predicted theoretically by Eq. (39), in particular, the angular dependences $h[\theta]$ and $g[\theta]$ stated in Eqs. (41) and (42), are shown as solid curves for the peaks centred around frequencies 00 (red), $f_{L}$ (green) and $2f_{L}$ (blue).
@@ IV Conclusions
On the one hand, the added noise amplifies the Larmor-induced peaks to be clearly visible within the power spectral density above the level of detection noise. As a result, our model describing the spin noise in the device can now be used by signal-processing tools interpret the detection data better—e.g., when operating the device as a scalar magnetometer, Bayesian inference methods such as the Kalman filter [48] can be applied to track fast changes of the strong magnetic field in real time beyond the magnetometer bandwidth [62, 63]. On the other hand, the induced noise can be used to naturally perform sensing tasks in the so-called covert manner [64], i.e. so that an adversary having access to the output of the magnetometer would not be able to recover the signal without possessing the pre-calibrated dynamical model that we propose. In this sense, we expect our model to be useful also for tracking time-varying signals encoded in oscillating RF-fields (amplitude/phase) directed perpendicularly to both the scalar field and the added noise.
@@
Experimental data created during this research are openly available from the University of Nottingham data repository at https://doi.org/10.17639/nott.7434.
@@ ACKNOWLEDGEMENTS
This work was supported by the QuantERA grant C’MON-QSENS! funded by the Engineering and Physical Sciences Research Council (EPSRC) (Grant No. EP/T027126/1). Project C’MON-QSENS! is also supported by the National Science Centre (2019/32/Z/ST2/00026), Poland under QuantERA, which has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement no. 731473. The work was also supported by the Novo Nordisk Foundation (Grant No. NNF20OC0064182) and the UK Quantum Technology Hub in Sensing and Timing, funded EPSRC (Grant No. EP/T001046/1) and EPSRC Grant No. EP/Y005260/1.
@@ Appendix A Evolution in the basis of spherical tensor operators
The evolution of the atomic ensemble without the effect of pumping, described by Eq. (19), is written in terms of the density matrix (1). In this appendix we show how to write it in terms of the vector of all spherical tensor components $\textbf{m}_{t}^{\mathrm{full}}$. In particular, since the right-hand side of the equation is linear in terms of $\rho$, we would like to find specific linear operators that act on $\textbf{m}_{t}^{\mathrm{full}}$ that correspond to specific operations performed on $\rho$, and enable us to write the dynamical equation in the form of (21). We also show that, since the resulting operators do not couple spherical tensor components of different rank, we only need the rank-2 component for a full description of the problem under consideration.
@@ A.1 Rotations induced by magnetic fields
Spherical-tensor operators, ${{\mathcal{T}}}^{(\kappa)}_{q}$, are defined by the decomposition of the density matrix (1) with respect to irreducible representations of the 3D-rotation group, i.e for each $\kappa=0,..,2F$, operators ${{\mathcal{T}}}^{(\kappa)}_{q}$ with $q=-\kappa,...,\kappa$ form a basis of a $(2\kappa+1)$-dimensional irreducible representation. These operators form a convenient basis for density matrices of a system with fixed angular momentum $F$:
@@
Using Eq. (51) we could analogously find that commuting the density matrix with $\hat{F}_{x}$ and $\hat{F}_{y}$ is equivalent to acting the operators $\textbf{J}^{(\kappa)}_{x}$ and $\textbf{J}^{(\kappa)}_{y}$, respectively, on $\textbf{m}_{t}^{(\kappa)}$. We use this to also obtain:
$\textbf{A}_{\mathrm{noise}}=-\mathrm{i}\sqrt{\omega_{\mathrm{noise}}}\textbf{J}^{(2)}_{x}.$ (53)
@@ B.1 Effective Lorentzian form
where $k_{\alpha}$, $k_{\beta}$ are the components of the k vector and $Q_{\alpha\beta}$ are the components of the matrix $\textbf{G}\,\bm{\sigma}\,\textbf{G}^{\mathrm{T}}$, in both cases written in the eigenbasis of F. As one can see, the resulting spectrum is a sum of symmetric (absorptive) and antisymmetric (dispersive) Lorentzian peaks, whose widths are the opposites of real parts ($\gamma_{\alpha}$), and central frequencies are the imaginary parts of the eigenvalues and their opposites (±$\pm\omega_{\alpha}$):
@@ Appendix C Analytic prediction of peak amplitudes
[…]
@@
because we need the second order of expansion in $\frac{1}{\Omega_{L}}$. Using this approximation we obtain Eqs. (39) from the main text, and:
@@ C.1 Irrelevance of the dispersive contributions
We note that we have not verified the predicted relations for the dispersive contributions of the line shape, see Eq. (91) and Eq. (92) in App. C. This is due to their contribution being below the noise floor of our system. We can verify that this is as expected for a Larmor frequency of approximately 9.45 kHz by calculating the ratios of the peak height equations and using the linewidths found in the experiment: $\Gamma_{1}=34$ Hz and $\Gamma_{2}=34$ Hz. We find $p^{d}_{1}/p^{a}_{1}$ from Eq. (91) divided by Eq. (39b) and we find $p^{d}_{2}/p^{a}_{2}$ from Eq. (92) divided by Eq. (39c). The calculations were done for the lowest white noise amplitude of $7~{}\mathrm{mV_{rms}}$ and the highest white noise amplitude of $3.5~{}\mathrm{V_{rms}}$. For $\mathrm{V_{noise}}=7~{}\mathrm{mV_{rms}}$ we find $p^{d}_{1}/p^{a}_{1}=1.1\times 10^{-4}$ and $p^{d}_{2}/p^{a}_{2}=3.2\times 10^{-4}$. Furthermore, when $\mathrm{V_{noise}}=3.5~{}\mathrm{V_{rms}}$ we find $p^{d}_{1}/p^{a}_{1}=3.1\times 10^{-5}$ and $p^{d}_{2}/p^{a}_{2}=1.8\times 10^{-4}$. Hence, the dispersive contribution is at least 4 orders of magnitude smaller than the absorptive contribution. Hence, this is significantly below our noise floor for both peaks, as seen experimentally.
Appendix D Amplitude calibration for the noise spectral density $\omega_{\mathrm{noise}}=2\pi f_{\mathrm{noise}}$
In order to determine the value of the constant $c$ relating the magnitude of the noise spectral density to the (RMS) amplitude of the voltage applied in the noise generator, i.e. $\omega_{\mathrm{noise}}=2\pi f_{\mathrm{noise}}=2\pi cV_{\mathrm{noise}}^{2}$, we directly measure the current in the coil that induces the noisy magnetic field. The signal from the generator, however, is first transformed thorough a 1 MHz low-pass filter to prevent the effect of aliasing, as well as a 1 kHz high-pass filter to eliminate any spurious contributions at very low (DC-like) frequencies. We record 100 traces of 0.01 s duration of the current reaching the coil (sampled at 40 MHz), while varying the voltage amplitude in the noise generator. Based on these we compute the PSD (30), which is presented in Fig. 7 in the units of $\mathrm{Hz}^{2}/\mathrm{Hz}$, after performing adequate rescaling given the values of the coil calibration factor (10.1 $\mathrm{nT/mV_{rms}}$) and the gyromagnetic ratio of caesium (3.5 Hz/nT).